OpenStudy (anonymous):

2. What is the simplified form of 8b^3c^2 + 4b^3c^2 ?

OpenStudy (anonymous):

I Guess all your Homework questions roaming over here :) Anyhow it would be 12b^3+6c^2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Collect like terms: \[8b^3c^2 + 4b^3c^2 = (8+4)b^3c^2 = 12b^3c^2\]

OpenStudy (anonymous):

Ohh i Just read out wrong..@@ Sorry but philips is right :)